OpenStudy (anonymous):

Let m1, r1 and v1 be the mean, the range and the variance of a group of numbers {x1,x2,x3...,x100} respectively. If m2, r2 and v2 are the mean, the range and the variance of the group of numbers {x1,x2,x3,...,x100,m1} respectively, which of the following must be true? I. m1=m2 II. r1=r2 III. v1=v2 i know the first one is correct, how about the others? @Callisto

OpenStudy (anonymous):

@Callisto i have done this question before but i forgot :O

OpenStudy (callisto):

If \(m_1\) is the mean of the first set of data, then \(x_1\le m_1 \le x_{100}\), agree?!

OpenStudy (anonymous):

why m1=x100 ??? :O

OpenStudy (anonymous):

second one i agree ... [?]

OpenStudy (callisto):

It may not be the same as x100, but if x1=x2=...=x100, then m1=x100

OpenStudy (callisto):

So, you agree that \(x_1\le m_1 \le x_{100}\)??

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Then adding \(m_1\) to the first data set won't affect the range since m1 does not change the extreme values (the max and the min.), agree?

OpenStudy (anonymous):

wait, let me organize o.0

OpenStudy (anonymous):

yup, i agree

OpenStudy (callisto):

So, 2) is true.

OpenStudy (anonymous):

what does 'variance' mean?

OpenStudy (callisto):

\(\sigma^2\), where \(\sigma\) = standard deviation.

OpenStudy (anonymous):

oh i see o.0

OpenStudy (callisto):

Do you know if (3) is true?

OpenStudy (anonymous):

the number of terms are not the same, so it's incorrect?

OpenStudy (callisto):

Hmm.. A better argument: For the first set of data, we have \(\sigma_1^2 = \frac{(x_1-m_1)^2+ .. + (x_{100}-m_1^2)}{100}\) For the second set of data, we have \(\sigma_2^2 = \frac{(x_1-m_1)^2+ .. + (x_{100}-m_1^2)+(m_1-m_1)^2}{101} = \frac{(x_1-m_1)^2+ .. + (x_{100}-m_1^2)}{101} <\sigma_1^2\) So, it is not true

OpenStudy (anonymous):

oh i forgot the standard deviation formula lol

OpenStudy (callisto):

Typo: It should be \((x_{100}-m_1)^2\) for all those three terms

OpenStudy (anonymous):

umm let me organize this first o.0

OpenStudy (callisto):

Take your time

OpenStudy (anonymous):

how do that two formula are out?

OpenStudy (anonymous):

whooops. oh i should clear the concepts on SD tonight lol thank you so muchhhhh <3

OpenStudy (callisto):

Welcome :)